# Local convergence rates for Wasserstein gradient flows and McKean-Vlasov equations with multiple stationary solutions

### v2

@@ Abstract
Non-linear versions of log-Sobolev inequalities, that link a free energy to its dissipation along the corresponding Wasserstein gradient flow (i.e. corresponds to Polyak-Lojasiewicz inequalities in this context), are known to provide global exponential long-time convergence to the free energy minimizers, and have been shown to hold in various contexts. However they cannot hold when the free energy admits critical points which are not global minimizers, which is for instance the case of the granular media equation in a double-well potential with quadratic attractive interaction at low temperature. This work addresses such cases, extending the general arguments when a log-Sobolev inequality only holds locally and, as an example, establishing such local inequalities for the granular media equation with quadratic interaction either in the one-dimensional symmetric double-well case or in higher dimension in the low temperature regime. The method provides quantitative convergence rates for initial conditions in a Wasserstein ball around the stationary solutions. The same analysis is carried out for the kinetic counterpart of the gradient flow, i.e. the corresponding Vlasov-Fokker-Planck equation. The local exponential convergence to stationary solutions for the mean-field equations, both elliptic and kinetic, is shown to induce for the corresponding particle systems a fast (i.e. uniform in the number or particles) decay of the particle system free energy toward the level of the non-linear limit.
@@ 1 Introduction
for some $\theta\in\mathbb{R}$. When $\theta\leqslant 0$ (repulsive interaction), or when $\theta>0$ and $\sigma^{2}/\theta$ is large enough (attractive interaction at high temperature or small interaction), there is a unique stationary solution to (1), which is globally attractive with exponential rate. However for a fixed $\theta>0$ there is a phase transition at some critical temperature $\sigma_{c}^{2}>0$ such that, for $\sigma^{2}<\sigma_{c}^{2}$, (1) with (2) admits three distinct stationary solutions (one symmetric, unstable, and two non-symmetric, stable) [46]. We are interested in this second case, and more precisely on obtaining local convergence rates, namely to prove that solutions which start close to a stable stationary solution converge exponentially fast to it. Among other motivations, this is an important question both for the theoretical understanding of metastable interacting particle systems, where the stability property of the non-linear limit drives the short-time behavior and induces the metastable behaviour [22, 12, 3, 20, 28, 5], and for practical questions of optimization in the Wasserstein space, for instance in the mean-field modelling of artificial neural networks [16, 29, 36, 35] (the loss function being convex only in toy models, like one-layer networks).
To illustrate the point, let us state a result obtained with our approach in the specific case of the symmetric double well (2). In the next statement (proven in Section 3.2.1, see Remarks 7 and 12) as in the rest of the work we write $\mathcal{W}_{2}$ and $\|\cdot\|_{TV}$ respectively the $L2$ Wasserstein distance and total variation norm and $\mathcal{P}_{2}(\mathbb{R}^{d})$ the set of probability measures on $\mathbb{R}^{d}$ with finite second moment.
@@
Notice that, in the sub-critical case, there is no assumption on $\sigma^{2}$ other than $\sigma^{2}<\sigma_{c}^{2}$, namely, the result holds arbitrarily close to the critical temperature (our results also apply in the simpler super-critical regime and provide, as soon as $\sigma^{2}>\sigma_{c}^{2}$, a global quantitative exponential convergence toward the unique stationary solution, see Remark 8). This is in contrast to the results of [49], also concerned with quantitative local convergence of (1), which require the temperature to be sufficiently small (for a given $\theta$). Moreover, the Wasserstein convergence in [49, Theorem 2.3] only gives a convergence speed of order $e^{-\lambda\sqrt{t}}$ for some $\lambda>0$, and additionally it requires $\theta>-\inf V^{\prime\prime}$, which our result does not (under this additional assumption we get that the rate $\lambda$ is uniform over $\sigma^{2}\in(0,\sigma_{0}^{2}]$ for any $\sigma_{0}^{2}\in(0,\sigma_{c}^{2})$, as in Proposition 16). The constants $\delta,\lambda,C$ in Proposition 1 can be made explicit (see Remark 9), namely the result is quantitative.
@@
$\partial_{t}\sqrt{f(x_{t})-f(x_{*})}\leqslant-\frac{1}{2\sqrt{\eta}}|\nabla f(x_{t})|\,,$
@@
We show that this argument extends to the infinite dimensional settings of gradient flows over the Wasserstein space (the lack of continuity of $f$ with respect to the Wasserstein distance in this case being circumvented by the regularization properties of (1) which, with the present notations, amounts to say that $x_{0}\mapsto f(x_{t})$ is continuous at $x_{*}$ as soon as $t>0$). The key new ingredient is thus the (local) dissipation inequality (3), which for elliptic McKean-Vlasov equations as (1) corresponds to a (local) non-linear log-Sobolev inequality (LSI). Such inequalities have been investigated in a number of works (see e.g. [11, 22, 26] and references within) but, to our knowledge, only in cases where they are global, corresponding in the finite-dimensional settings above to the case where (3) holds for all $x\in\mathbb{R}^{d}$ (for instance when $f$ is uniformly strongly convex). Our main contribution is thus to show that the method also works locally and, more importantly, to show that it is indeed possible to establish such local dissipation inequalities in some cases where the global inequality fails. More specifically, the main results of this work are the following:
@@
Last, let us mention that for McKean-Vlasov equations which do not necessarily write as Wasserstein gradient flows, but under different assumptions than ours, a similar statement to Proposition 1 was recently obtained in [19]. It provides exponential convergence when the initial condition is in a small $\mathcal{W}_{1}$ neighborhood of a stationary solution. The method is completely different and relies on the differentiation, in the sense of Lions derivatives, of the drift of the underlying nonlinear SDE in the neighborhood of the considered invariant measure, which then provides a criterion for the stability of the invariant measure.
@@ 2.1 General settings, assumptions and notations
for which $E_{\rho}(x)=V(x)+\rho\star W(x)$ and Assumptions 1 and 2 are satisfied as soon as, for instance, $V$ is convex outside a compact set and $W$ is lower bounded with at most a quadratic growth at infinity (other conditions can be considered, for instance in repulsive cases $W$ may not be bounded below but Assumption 1 still holds if $V$ grows faster than $-W$ at infinity) Under suitable regularity condition, a straightforward computation gives
$\partial_{t}\mathcal{F}(\rho_{t})=-\int_{\mathbb{R}^{d}}\left|\nabla\frac{\delta\mathcal{F}(\rho_{t})}{\delta\mu}\right|^{2}\mathrm{d}\rho_{t}\,.$ (11)
In particular, the free energy is decreasing along time, and the stationary solutions $\rho_{*}$ are the critical points of $\mathcal{F}$, characterized by the fact that $\frac{\delta\mathcal{F}}{\delta\mu}(\rho_{*})$ is constant, which, using that $\frac{\delta\mathcal{H}}{\delta\mu}(\rho)=\ln\rho$, is equivalent to the self-consistency equation
@@
$\partial\tilde{\rho}_{t}=\sigma^{2}\Delta\tilde{\rho}_{t}+\nabla\cdot\left(\tilde{\rho}_{t}\nabla E_{\rho}\right)\,,$
@@
$\mathcal{I}(\nu|\mu)=\int_{\mathbb{R}^{d}}\left|\nabla\ln\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{2}\mathrm{d}\nu\,.$
@@ 2.2 Two known cases with global convergence
The convexity condition (14) is known to hold in various settings, in particular for mean-field models of one-layer neuron networks [29, 16]. In the granular media case (10), assume that the interaction potential is of the form
@@
so that (14) holds for all $\mu_{0},\mu_{1},t$ if and only if $a_{k}=0$ for all $k\in\mathbb{N}$, which corresponds to repulsive interaction (and for quadratic potentials (2) to the case where $\theta\leqslant 0$, i.e. $W$ is concave). See also [15, Section 3.1] for further examples satisfying (14).
@@
$\partial_{t}\mathcal{H}(\rho_{t}|\rho_{*})=-\sigma^{2}\mathcal{I}(\rho_{t}|\rho_{*})+\int_{\mathbb{R}^{d}}\left(\nabla E_{\rho_{*}}-\nabla E_{\rho_{t}}\right)\cdot\nabla\ln\left(\frac{\rho_{t}}{\rho_{*}}\right)\mathrm{d}\rho_{t}\,.$
@@
$\partial_{t}\mathcal{H}(\rho_{t}|\rho_{*})$$\leqslant$ $-\sigma^{2}\mathcal{I}(\rho_{t}|\rho_{*})+\left\|\nabla E_{\rho_{*}}-\nabla E_{\rho_{t}}\right\|_{\infty}\sqrt{\mathcal{I}(\rho_{t}|\rho_{*})}$
@@
$\leqslant$ $\left(-\frac{\sigma^{2}}{2\eta}+\frac{2L^{2}\eta}{\sigma^{2}}\right)\mathcal{H}(\rho_{t}|\rho_{*})\,,$
@@ 2.3 Lower-bounded curvature
and thus, interverting the roles of $\mu_{1}$ and $\mu_{0}$,
@@ 2.4 Local non-linear log-Sobolev inequalities
The first inequality is straightforward from (11) and (NL-LSI). The proof of the second inequality is the same as the proof of [22, Theorem 3.2] (although a factor 2 seems to disappear wrongly in the latter when using the Benamou-Brenier formula, which explains the difference with our result ; notice that our factor 4 is consistent with the classical linear case [41], or the finite-dimensional settings (5)). Since (9) can be written as the continuity equation
@@
$\partial_{t}\mathcal{F}_{\mathcal{A}}(\rho_{t})=-\sigma^{4}\mathcal{I}\left(\rho_{t}|\Gamma(\rho_{t})\right)\\ \leqslant-\sqrt{\sigma^{4}\mathcal{I}\left(\rho_{t}|\Gamma(\rho_{t})\right)\mathcal{F}_{\mathcal{A}}(\rho_{t})/\overline{\eta}}\,,$
@@
Rearranging the terms and using that $\sqrt{\mathcal{F}_{\mathcal{A}}(\rho_{t})}\geqslant 0$ concludes. ∎
As a consequence, when (NL-LSI) holds in a set $\mathcal{A}\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ that contains a unique critical point $\rho_{*}\in\mathcal{K}$, in order to get the long-time convergence to $\rho_{*}$ at exponential speed starting from an initial condition $\rho_{0}\in\mathcal{A}$, it mainly remains to show that $T_{\mathcal{A}}(\rho_{0})=+\infty$. In specific cases, this can be shown by various means, for instance in dimension 1 by monotonicity arguments as in [48], or in general using that the free energy decreases as in [47], etc. In the following, we focus on a self-contained argument which establishes the exponential convergence towards $\rho_{*}$ under the natural condition that $\mathcal{W}_{2}(\rho_{*},\rho_{0})$ is sufficiently small (in particular, without assuming that $\mathcal{F}(\rho_{0})<\infty$).

### v4

@@ Abstract
Non-linear versions of log-Sobolev inequalities, that link a free energy to its dissipation along the corresponding Wasserstein gradient flow (i.e. corresponds to Polyak-Lojasiewicz inequalities in this context), are known to provide global exponential long-time convergence to the free energy minimizers, and have been shown to hold in various contexts. However they cannot hold when the free energy admits critical points which are not global minimizers, which is for instance the case of the granular media equation in a double-well potential with quadratic attractive interaction at low temperature. This work addresses such cases, extending the general arguments when a log-Sobolev inequality only holds locally and, as an example, establishing such local inequalities for the granular media equation with quadratic interaction either in the one-dimensional symmetric double-well case or in higher dimension in the low temperature regime. The method provides quantitative convergence rates for initial conditions in a Wasserstein ball around the stationary solutions. The same analysis is carried out for the kinetic counterpart of the gradient flow, i.e. the corresponding Vlasov-Fokker-Planck equation. The local exponential convergence to stationary solutions for the mean-field equations, both elliptic and kinetic, is shown to induce for the corresponding particle systems a fast (i.e. uniform in the number of particles) decay of the particle system free energy toward the level of the non-linear limit.
@@ 1.1 Motivation
for some $\theta\in\mathbb{R}$. When $\theta\leqslant 0$ (repulsive interaction), or when $\theta>0$ and $\sigma^{2}/\theta$ is large enough (attractive interaction at high temperature or small interaction), there is a unique stationary solution to (1), which is globally attractive with exponential rate. However for a fixed $\theta>0$ there is a phase transition at some critical temperature $\sigma_{c}^{2}>0$ such that, for $\sigma^{2}<\sigma_{c}^{2}$, (1) with (2) admits three distinct stationary solutions (one symmetric, unstable, and two non-symmetric, stable) [50]. We are interested in this second case, and more precisely on obtaining local convergence rates, namely to prove that solutions which start close to a stable stationary solution converge exponentially fast to it. Among other motivations, this is an important question both for the theoretical understanding of metastable interacting particle systems, where the stability property of the non-linear limit drives the short-time behavior and induces the metastable behavior [23, 13, 3, 21, 30, 5], and for practical questions of optimization in the Wasserstein space, for instance in the mean-field modelling of artificial neural networks [17, 31, 38, 37] (the loss function being convex only in toy models, like one-layer networks, cf. (21)).
To illustrate the point, let us state a result obtained with our approach in the specific case of the symmetric double well (2). In the next statement (proven in Section 3.2.1, see Remarks 8 and 13) as in the rest of the work we write $\mathcal{W}_{2}$ and $\|\cdot\|_{TV}$ respectively the $L2$ Wasserstein distance and total variation norm and $\mathcal{P}_{2}(\mathbb{R}^{d})$ the set of probability measures on $\mathbb{R}^{d}$ with finite second moment.
@@
Notice that, in the sub-critical case, there is no assumption on $\sigma^{2}$ other than $\sigma^{2}<\sigma_{c}^{2}$, namely, the result holds arbitrarily close to the critical temperature (our results also apply in the simpler super-critical regime and provide, as soon as $\sigma^{2}>\sigma_{c}^{2}$, a global quantitative exponential convergence toward the unique stationary solution, see Remark 9). This is in contrast to the results of [53], also concerned with quantitative local convergence of (1), which require the temperature to be sufficiently small (for a given $\theta$). Moreover, the Wasserstein convergence in [53, Theorem 2.3] only gives a convergence speed of order $e^{-\lambda\sqrt{t}}$ for some $\lambda>0$, and additionally it requires $\theta>-\inf V^{\prime\prime}$, which our result does not (under this additional assumption we get that the rate $\lambda$ is uniform over $\sigma^{2}\in(0,\sigma_{0}^{2}]$ for any $\sigma_{0}^{2}\in(0,\sigma_{c}^{2})$, as in Proposition 16). The constants $\delta,\lambda,C$ in Proposition 1 can be made explicit (see Remark 10), namely the result is quantitative.
1.2 The strategy in finite dimension
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\sqrt{f(x_{t})-f(x_{*})}\leqslant-\frac{1}{2\sqrt{\eta}}|\nabla f(x_{t})|\,,$ (5)
@@ 1.3 Extension to the Wasserstein space
We show that the previous argument extends to the infinite dimensional settings of gradient flows over the Wasserstein space (the lack of continuity of $f$ with respect to the Wasserstein distance in this case being circumvented by the regularization properties of (1) which, with the present notations, amounts to say that $x_{0}\mapsto f(x_{t})$ is continuous at $x_{*}$ as soon as $t>0$). The key new ingredient is thus the (local) dissipation inequality (3), which for elliptic McKean-Vlasov equations as (1) corresponds to a (local) non-linear log-Sobolev inequality (LSI). Such inequalities have been investigated in a number of works (see e.g. [12, 23, 28] and references within) but, to our knowledge, only in cases where they are global, corresponding in the finite-dimensional settings above to the case where (3) holds for all $x\in\mathbb{R}^{d}$ (for instance when $f$ is uniformly strongly convex). Our main contribution is thus to show that the method also works locally and, more importantly, to show that it is indeed possible to establish such local dissipation inequalities in some cases where the global inequality fails.
As an illustration in a specific case, let us sum up the main steps of the proof of Proposition 1.
  1. 1.
    The first step is to prove the analogue of (3) in the present context. Since the dissipation of the free energy along the flow reads
    $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}(\rho_{t})=-\sigma^{4}\mathcal{I}\left(\rho_{t}|\Gamma(\rho_{t})\right)$
    with $\mathcal{I}$ the Fisher information and $\Gamma(\rho)$ the local equilibrium (see (18) and (17)), the goal is to find a constant $\eta>0$ and a set of probability measures $\mathcal{A}$ such that
    $\forall\rho\in\mathcal{A},\qquad\mathcal{F}(\rho)-\inf_{\mathcal{A}}\mathcal{F}\leqslant\eta\mathcal{I}\left(\rho|\Gamma(\rho)\right)\,.$ (7)
    To do so, we notice that $\mathcal{F}$ can be decomposed as
    $\mathcal{F}(\rho)=\mathcal{H}\left(\rho|\Gamma(\rho)\right)+g\left(m(\rho)\right)$
    for some function $g:\mathbb{R}\mapsto\mathbb{R}$ with the mean $m(\rho)=\int_{\mathbb{R}}x\rho(x)\mathrm{d}x$, where $\mathcal{H}$ stands for the relative entropy (cf. Section 3.1.2). Classical arguments show that $\Gamma(\rho)$ satisfies a so-called log-Sobolev inequality, with a constant independent from $\rho$, which shows that $\mathcal{H}(\rho|\Gamma(\rho))\leqslant\eta^{\prime}\mathcal{I}(\rho|\Gamma(\rho))$ for all $\rho$ for some $\eta^{\prime}>0$. Hence, it only remains to analyze a $1d$ function $g$. In the context of Proposition 1, we can actually bound
    $\mathcal{F}(\rho)-\inf\mathcal{F}=\mathcal{F}(\rho)-\mathcal{F}(\rho_{*})=\mathcal{H}(\rho|\Gamma(\rho)+g(m(\rho))-g(m(\rho_{*}))\\ \leqslant\mathcal{H}(\rho|\Gamma(\rho))+C|m(\rho)-m(\rho_{*})|^{2}$
    for some $C>0$. The main step is thus to analyze a $1d$ fixed-point problem associated to $g$ leading to a bound of the form
    $|m(\rho)-m(\rho_{*})|\leqslant C_{\varepsilon}|m(\rho)-m\left(\Gamma(\rho)\right)|$ (8)
    for some $C_{\varepsilon}>0$, for all $\rho$ with $|m(\rho)|\geqslant\varepsilon>0$, for all $\varepsilon>0$ (see Proposition 13). Since $|m(\rho)-m(\Gamma(\rho))|\leqslant\mathcal{W}_{2}(\rho,\Gamma(\rho))|\leqslant 2\eta^{\prime}\sqrt{\mathcal{I}(\rho|\Gamma(\rho))}$ thanks to the log-Sobolev inequality satisfied by $\Gamma(\rho)$ (cf. the proof of Proposition 9 for details), this concludes the proof of (7) over the set $\mathcal{A}=\{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d}),|m(\rho)|\geqslant\varepsilon\}$, for an $\varepsilon>0$ smaller than $|m(\rho_{*})|$. In the criticial case where $\sigma=\sigma_{c}$ , the fixed point problem associated to $g$ is degenerate and the bound (8) has to be replaced by
    $|m(\rho)-m(\rho_{*})|\leqslant C_{\varepsilon}\left(|m(\rho)-m\left(\Gamma(\rho)\right)|+|m(\rho)-m\left(\Gamma(\rho)\right)|^{1/3}\right)\,,$
    cf. Proposition 15.
  2. 2.
    The Benamou-Brenier formulation of the Wasserstein distance gives the analogue of (6), namely
    $\mathcal{W}_{2}^{2}(\rho_{t},\rho_{0})\leqslant C\left(\mathcal{F}(\rho_{0})-\mathcal{F}(\rho_{*})\right)$ (9)
    for some $C>0$, as long as $\rho_{t}\in\mathcal{A}$ (cf. Lemma 4).
  3. 3.
    Thanks to the regularization property of the elliptic PDE (1), a Wang-Harnak inequality holds : for some $C>0$,
    $\mathcal{F}(\rho_{1})-\mathcal{F}(\rho_{*})\leqslant C\mathcal{W}_{2}^{2}(\rho_{*},\rho_{0})\,,$
    cf. Corollary 7. Combining this with (9) and a simple stability bound $\mathcal{W}_{2}(\rho_{t},\rho_{*})\leqslant C\mathcal{W}_{2}(\rho_{0},\rho_{*})$ for some $C>0$ for $t\in[0,1]$ gives
    $\mathcal{W}_{2}^{2}(\rho_{t},\rho_{0})\leqslant C^{\prime}\mathcal{W}_{2}^{2}(\rho_{*},\rho_{0})\,,$
    for some $C^{\prime}>0$, as long as as $\rho_{t}\in\mathcal{A}$. This allows to conclude as in the finite dimensional case with (6), despite $\mathcal{F}$ not being continuous with respect to $\mathcal{W}_{2}$. In particular, since $\mathcal{A}$ contains a Wasserstein ball centered on $\rho_{*}$, reasoning as in finite dimension, we get that there exists a smaller ball from which solutions started there remain in $\mathcal{A}$ for all times. Conclusion follows as in the finite dimensional case.
1.4 Main contributions and organization
To conclude this introduction, let us describe the main results of this work.
@@ 1.4 Main contributions and organization
Last, let us mention that for McKean-Vlasov equations which do not necessarily write as Wasserstein gradient flows, but under different assumptions than ours, a similar statement to Proposition 1 was recently obtained in [20]. It provides exponential convergence when the initial condition is in a small $\mathcal{W}_{1}$ neighborhood of a stationary solution. The method is completely different and relies on the differentiation, in the sense of Lions derivatives, of the drift of the underlying nonlinear SDE in the neighborhood of the considered invariant measure, which then provides a criterion for the stability of the invariant measure. In the spirit of earlier works such as [48], it doesn’t give very explicit convergence rates or stability radius.
@@ 2.1 General settings, assumptions and notations
for which $E_{\rho}(x)=V(x)+\rho\star W(x)$ and Assumptions 1 and 2 are satisfied as soon as, for instance, $V$ is convex outside a compact set and $W$ is lower bounded with at most a quadratic growth at infinity (other conditions can be considered, for instance in repulsive cases $W$ may not be bounded below but Assumption 1 still holds if $V$ grows faster than $-W$ at infinity) Under suitable regularity conditions, a straightforward computation gives
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}(\rho_{t})=-\int_{\mathbb{R}^{d}}\left|\nabla\frac{\delta\mathcal{F}(\rho_{t})}{\delta\mu}\right|^{2}\mathrm{d}\rho_{t}\,,$ (15)
see e.g. [12, Proposition 2.1]. In particular, the free energy is decreasing along time, and the stationary solutions $\rho_{*}$ are the critical points of $\mathcal{F}$, characterized by the fact that $\frac{\delta\mathcal{F}}{\delta\mu}(\rho_{*})$ is constant, which, using that $\frac{\delta\mathcal{H}}{\delta\mu}(\rho)=\ln\rho$, is equivalent to the self-consistency equation
@@
$\partial_{t}\tilde{\rho}_{t}=\sigma^{2}\Delta\tilde{\rho}_{t}+\nabla\cdot\left(\tilde{\rho}_{t}\nabla E_{\rho}\right)\,,$
@@
$\mathcal{I}(\nu|\mu)=\int_{\mathbb{R}^{d}}\left|\nabla\ln\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{2}\mathrm{d}\nu\,.$ (18)
@@ 2.2 Two known cases with global convergence
The convexity condition (19) is known to hold in various settings, in particular for mean-field models of one-layer neuron networks [31, 17], where
$\mathcal{E}\left(\mu\right)=\int_{\mathbb{R}^{n+m}}\left|z-\int_{\mathbb{R}^{d}}\Phi(x,y)\mu(\mathrm{d}x)\right|^{2}m(\mathrm{d}y\mathrm{d}z)+c\int_{\mathbb{R}^{d}}|x|^{2}\mu(\mathrm{d}x)\,,$ (21)
where $m$ is the distribution of the data ($y\in\mathbb{R}^{n}$ being the input, $z\in\mathbb{R}^{m}$ the output), $x$ is the parameter of a single neuron, $\Phi(x,\cdot)$ is the map parametrized by a neuron with parameters $x$, $\mu$ is the distribution of the parameters in the neurons of the layer, and $c>0$ is a penalization parameter. In the granular media case (14), assume that the interaction potential is of the form
@@
so that (19) holds for all $\mu_{0},\mu_{1},t$ if and only if $a_{k}=0$ for all $k\in\mathbb{N}$, which corresponds to repulsive interaction (and for quadratic potentials (2) to the case where $\theta\leqslant 0$, i.e. $W$ is concave). See also [16, Section 3.1] for further examples satisfying (19).
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(\rho_{t}|\rho_{*})=-\sigma^{2}\mathcal{I}(\rho_{t}|\rho_{*})+\int_{\mathbb{R}^{d}}\left(\nabla E_{\rho_{*}}-\nabla E_{\rho_{t}}\right)\cdot\nabla\ln\left(\frac{\rho_{t}}{\rho_{*}}\right)\mathrm{d}\rho_{t}\,.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(\rho_{t}|\rho_{*})$$\leqslant$ $-\sigma^{2}\mathcal{I}(\rho_{t}|\rho_{*})+\left\|\nabla E_{\rho_{*}}-\nabla E_{\rho_{t}}\right\|_{\infty}\sqrt{\mathcal{I}(\rho_{t}|\rho_{*})}$
@@
$\leqslant$ $\left(-\frac{\sigma^{2}}{2\eta}+\frac{2L^{2}\eta}{\sigma^{2}}\right)\mathcal{H}(\rho_{t}|\rho_{*})\,.$
@@ 2.3 Lower-bounded curvature
and thus, switching the roles of $\mu_{1}$ and $\mu_{0}$,
@@ 2.4 Local non-linear log-Sobolev inequalities
The first inequality is straightforward from (15) and (NL-LSI). The proof of the second inequality is the same as the proof of [23, Theorem 3.2] (although a factor 2 seems to disappear wrongly in the latter when using the Benamou-Brenier formula, which explains the difference with our result; notice that our factor 4 is consistent with the classical linear case [44], or the finite-dimensional settings (6)), as we now detail. Since (13) can be written as the continuity equation
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}_{\mathcal{A}}(\rho_{t})=-\sigma^{4}\mathcal{I}\left(\rho_{t}|\Gamma(\rho_{t})\right)\\ \leqslant-\sqrt{\sigma^{4}\mathcal{I}\left(\rho_{t}|\Gamma(\rho_{t})\right)\mathcal{F}_{\mathcal{A}}(\rho_{t})/\overline{\eta}}\,,$
@@
Rearranging the terms and using that $\sqrt{\mathcal{F}_{\mathcal{A}}(\rho_{t})}\geqslant 0$ concludes the proof of (32). ∎
As a consequence, when (NL-LSI) holds in a set $\mathcal{A}\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ that contains a unique a local minimizer $\rho_{*}\in\mathcal{K}$ and no other critical point, in order to get the long-time convergence to $\rho_{*}$ at exponential speed starting from an initial condition $\rho_{0}\in\mathcal{A}$, it mainly remains to show that $T_{\mathcal{A}}(\rho_{0})=+\infty$. In specific cases, this can be shown by various means, for instance in dimension 1 by monotonicity arguments as in [52], or in general using that the free energy decreases as in [51], etc. In the following, we focus on a self-contained argument which establishes the exponential convergence towards $\rho_{*}$ under the natural condition that $\mathcal{W}_{2}(\rho_{*},\rho_{0})$ is sufficiently small (in particular, without assuming that $\mathcal{F}(\rho_{0})<\infty$).